# Microphysical Regulation of Non-Ideal MHD in Weakly-Ionized Systems: Does the Hall Effect Matter?

### v1

@@ Acknowledgements.
Appendix A Additional Multi-Species Details
In the main text, for the arbitrary multi-species case (§ 2.4.2), we propose a simple weighting of the effective Hall/Ohmic/ambipolar coefficients to capture the leading-order anomalous resistivity effects. Here we discuss the standard multi-species derivation of the non-ideal terms in more detail, to show that our proposed corrects capture the relevant behavior in the limits of interest.
@@ A.1 Overview and Classical Derivation
In principle, one could take this approach and simply use the drift/slip-velocity-dependent form of all the collision rates $\omega_{ij}\rightarrow\omega_{ij}({\bf V},\,...)+\omega^{\rm an}_{ij}({\bf V},\,...)$ (and similarly modify the inelastic and charge exchange rates $\dot{\rho}$ to be appropriate functions of ${\bf V}$ as well), i.e. account for both the classical dependence of the direct collision rates $\omega_{ij}({\bf V},\,...)$ on all the relative velocities ${\bf u}_{j}$ (in ${\bf V}$; e.g. Pinto & Galli 2008) as well as the anomalous resistivity/collision rate of each species with respect to others (itself also a function of ${\bf V}$ through e.g. functions like $\Theta({\bf u}_{j}-{\bf u}_{i})$ in Eq. 16). But then the terms appearing in ${\bf S}$ are highly non-linear functions of the different components of ${\bf V}$ and (implicitly) ${\bf E}$ and ${\bf J}$, and multi-dimensional iterative root-finding methods are needed. And even if we took the “zero drift/slip” coefficients ($\omega_{ij}({\bf V})\rightarrow\omega_{ij}({\bf V}=\boldsymbol{0})$) so ${\bf S}$ were independent of ${\bf V}$, if we retain general cross-interaction terms $\omega_{ij}$ and charge-exchange/inelastic terms then the matrix solve remains quite complicated. Still, for methods that actually numerically solve a chemical network on-the-fly for the abundances of $N$ directly, this is not necessarily more expensive than the network solve itself, so may be preferred. However, most implementations either pre-compute chemical tables or otherwise simplify to only use some end product of this series of inversions. Moreover, even if we followed the non-linear procedure above, for general $\omega_{ij}({\bf V})$ and $N$, there is no guarantee of a unique solution ${\bf V}$ for given ${\bf J}$ (there can be multiple viable sets of ${\bf u}_{j}$, which would require evolving $D_{t}{\bf u}_{j}$ to distinguish).
@@
We can now see the primary motivation for the common literature assumptions (3) (assume dominant neutrals with ${\bf U}\approx{\bf u}_{n}$, and neglect all $\dot{\rho}_{ij}$ and non-neutral collision $\omega_{ij}$ terms) and (4) (assume vanishing drift/slip so $\omega_{ij}\rightarrow\omega_{ij}({\bf V}=\boldsymbol{0})$). With these simplifications, ${\bf S}$ has a trivial block form and the equation for each species is separable: ${\bf W}\equiv(c/B)\,({\bf E}+{\bf U}\times{\bf B}/c)\approx{\bf b}\times\delta{\bf u}_{jn}+\tilde{\beta}_{j}^{-1}\delta{\bf u}_{jn}$ (where $\delta{\bf u}_{jn}\equiv{\bf u}_{j}-{\bf u}_{n}$ are the slip velocities and $\tilde{\beta}_{j}\equiv q_{j}B/m_{j}c\omega_{jn}$). This gives the familiar result that the relative contributions to the current for each species depend only on $\tilde{\beta}_{j}$, and coefficients $\eta_{O}=(cB/4\pi)\tilde{\sigma}_{O}^{-1}$, $\eta_{H}=(cB/4\pi)\tilde{\sigma}_{H}/\tilde{\sigma}_{\bot}^{2}$, $\eta_{A}=(cB/4\pi)(\tilde{\sigma}_{P}/\tilde{\sigma}_{\bot}^{2}-1/\tilde{\sigma}_{O})$ where $\tilde{\sigma}_{O}\equiv\sum n_{j}q_{j}\tilde{\beta}_{j}$, $\tilde{\sigma}_{H}\equiv n_{j}q_{j}/(1+\tilde{\beta}_{j}^{2})$, $\tilde{\sigma}_{P}\equiv n_{j}q_{j}\tilde{\beta}_{j}/(1+\tilde{\beta}_{j}^{2})$, $\tilde{\sigma}_{\bot}^{2}\equiv\tilde{\sigma}_{H}^{2}+\tilde{\sigma}_{P}^{2}$. With the slip velocities [^7] The slip velocities are given in a frame where ${\bf b}=\{0,0,1\}=\hat{z}$ by: $\delta{\bf u}_{jn,x}$ $=\frac{\tilde{\beta}_{j}(W_{x}+\tilde{\beta}_{j}W_{y})}{(1+\tilde{\beta}_{j}^{2})}=\frac{\tilde{\beta}_{j}[J_{x}(\tilde{\sigma}_{P}-\tilde{\beta}_{j}\tilde{\sigma}_{H})+J_{y}(\tilde{\sigma}_{H}+\tilde{\beta}_{j}\tilde{\sigma}_{P})]}{(1+\tilde{\beta}_{j}^{2})\sigma_{\bot}^{2}}\ ,$ $\delta{\bf u}_{jn,y}$ $=\frac{\tilde{\beta}_{j}(W_{y}-\tilde{\beta}_{j}W_{x})}{(1+\tilde{\beta}_{j}^{2})}=\frac{\tilde{\beta}_{j}[J_{y}(\tilde{\sigma}_{P}-\tilde{\beta}_{j}\tilde{\sigma}_{H})-J_{x}(\tilde{\sigma}_{H}+\tilde{\beta}_{j}\tilde{\sigma}_{P})]}{(1+\tilde{\beta}_{j}^{2})\sigma_{\bot}^{2}}\ ,$ $\delta{\bf u}_{jn,z}$ $=\tilde{\beta}_{j}W_{z}=\frac{\tilde{\beta}_{j}J_{z}}{\tilde{\sigma}_{O}}\ .$ (24) $\delta{\bf u}_{j}$, we can calculate each pairwise drift ($\delta{\bf u}_{j}-\delta{\bf u}_{i}$), and again in principle modify $\omega_{ij}$ to account for each drift+slip speed. But this again introduces the difficulties above: we then modify $\tilde{\beta}_{j}$ so must iteratively re-calculate the relative drift/slip speeds, are not ensured a unique solution, and (conceptually) must evaluate “which drifts matter.” [^8] For example, even if electrons carry most of the current, since we only retain their effective collision rate with neutrals here, do we include anomalous resistivity if their drift is subthermal relative to a dominant ion in the current but superthermal relative to a more trace ion, like charged grains?
@@
Figure 4: Ohmic, Hall, and ambipolar diffusivities $\eta_{O,H,A}$ as a function of density as in Fig. 2 (for the “circumstellar disk/MMSN” parameters therein), but for the specific numerical cases in § A.3. Left: Comparison of a generally-subthermal case ($\ell_{B}=10\,\ell_{\rm crit}$; dotted) to a superthermal drift case ($\ell_{B}=0.1\,\ell_{\rm crit}^{\rm drift}$; solid). Right: Same but for the superthermal slip case $\ell_{B}=0.1\,\ell_{\rm crit}^{\rm slip}$. For each, we compare (1) the prediction from the simple parameterized rescaling we propose in Eq. 16 (for the multi-species scalings in § 2.4.2; “Approximate” in thin lines); to (2) the case where ignore charged dust grains as a current carrier (treat them as part of the neutral inertia taking $\tilde{\beta}_{\rm grain}\rightarrow 0$) and solve the full non-linear three-species problem with the exact coefficients (§ 2.1; “Exact” 3-species in intermediate lines); and (2) the results of exactly solving the non-linear multi-species equations as described in § A.1 with all possible pairs of species featuring simple Epstein-type scalings and anomalous terms (§ A.3; “Exact” 5-species in thick lines). The 3 and 5-species methods give excellent agreement, indicating that grain current and differences between exact velocity-dependent scalings in the superthermal regime make little difference. The closed-form analytic approximate scaling from Eq. 16 reproduces these reasonably well in all regimes of interest. Key qualitative results (e.g. vanishing of the Hall-dominated regime) are robust.
@@ A.2 Application to Scalings In the Text
A.3 A Specific Example
@@
To numerically validate the above and our approach, we revisit our simple example in the text of Fig. 2, where we calculate the $\eta_{O,H,A}$ coefficients according to the simple five-species model from § 2.4.3, but solving the generalized non-linear multidimensional problem exactly as described above (§ A.1; making assumptions (1), (2), (5) to close the system, but no others). We now have interactions between electrons, heavy ions, neutrals, positively and negatively charged grains. Per § A.1, to even write down the equations for ${\bf E}$ in this limit we must define the velocity-dependent interaction terms between all pairs of particles, each of which can have different instability thresholds and qualitative behaviors of their anomalous terms (many of which remain poorly understood in detail). So we necessarily adopt a simple toy model, with the purpose not of providing any sort of “first principles” calculation, but simply to validate the simple approximate scalings discussed above. We neglect charge-exchange reactions, but consider all effective collisional cross terms. This includes both a “classical” (but correctly velocity-dependent) rate which we approximate with a simple hard-sphere-scattering type expression $\omega^{\rm coll}_{ji}=\omega^{\rm coll}_{ji}(\delta{\bf u}_{ji}\rightarrow 0)\times\left[1+|\delta{\bf u}_{ji}|^{2}/v_{T,\,ij}^{2}\right]^{1/2}$ (for which we use the values of $\langle\sigma v(\delta{\bf u}_{ji}\rightarrow 0)\rangle$ in text, supplemented with the zero-drift ion-grain and electron-grain cross sections from Draine & Sutin 1987), plus a pairwise anomalous term to capture different instabilities that can occur between each pair of interacting charged species, $\omega^{\rm an}_{ji}={\rm MAX}[\Omega_{j},\,\Omega_{i},\,\omega_{pj},\,\omega_{pi}]\,\Theta(|\delta{\bf u}_{ji}|/v_{T,ij})$ for $j,\,i\neq n$, with $v_{T,\,ij}={\rm MIN}[v_{T,\,i},\,v_{T,\,j}]$ (per Eqs. 20-22, with $\psi_{j}=0$ for $j\neq i,n$, so just the effective speeds of the individual carriers plus potentially-entrained neutrals for each). Like in Fig. 2 we consider two limits, which we define by the magnetic-field gradient $\ell_{B}=10\,\ell_{\rm crit}$ and $\ell_{B}=0.1\,\ell_{\rm crit}^{\rm drift}$ or $\ell_{B}=0.1\,\ell_{\rm crit}^{\rm slip}$ (with $v_{T}={\rm MIN}[v_{T,\,n},\,v_{A,\,{\rm ideal}}]$ in $\ell_{\rm crit}$). In the model in the text, these correspond one-to-one to the sub-thermal ($v_{\rm drift}\sim 0.1\,v_{T}$) and super-thermal ($v_{\rm drift}\sim 10\,v_{T}$) limits, but since there is no single drift velocity to speak of in the more general case, they must be defined in terms of $\ell_{B}$ (which is independent of the collision rates and species involved). Because of the non-linear dependencies involved, we must also specify the geometry, namely $\cos{\theta}\equiv\hat{\bf J}\cdot{\bf b}$, for which we take a representative $\theta=30^{\circ}$, but note the results are very weakly dependent on $\theta$, except if very close to parallel ($|\cos{\theta}|\approx 1$) where the ambipolar and Hall terms trivially vanish. Again, we emphasize the various coefficients and rates above are not meant to be exact, but simply to provide some calculable illustration.
@@
We further consider one more comparison. If we treat the dust grains as simply a “charge sink” and assume their mass to be anchored to the neutrals (adding to neutral inertia to leading order), then our system reduces to a three-component (neutrals, ions, electrons) system. We take this, plus the full velocity-dependent scattering rates from Pinto & Galli (2008) for each component, with a single ion-electron anomalous resistivity added to the ion-electron collision rate (where present), and otherwise adopt the scalings above. Then we can apply our solution from the main text (§ 2.1) and exactly solve the non-linear problem with all terms from § 2.1 included.

### v2

@@ Acknowledgements.
Appendix A Additional Multi-Species Details, Numerical Solutions, and Comparisons
In the main text, for the arbitrary multi-species case (§ 2.4.2), we propose a simple weighting of the effective Hall/Ohmic/ambipolar coefficients to capture the leading-order anomalous resistivity effects, together with the approximate analytic rescalings in Eq. 16. Here we discuss the standard multi-species derivation of the non-ideal terms in more detail, and calculate more accurate solutions numerically, to show that our proposed corrections capture the relevant behavior in the limits of interest.
@@ A.1 Overview and Classical Derivation
[…]
However, most implementations either pre-compute chemical tables or otherwise simplify to only use some end product of this series of inversions. Moreover, even if we followed the non-linear procedure above, for general $\omega_{ij}({\bf V})$ and $N$, there is no guarantee of a unique solution $\bf V$} for given $\bf J$} (there can be multiple viable sets of ${\bf u}_{j}$, which would require evolving $D_{t}{\bf u}_{j}$ to distinguish, depending on how the $\omega_{ij}$ terms depend on $\bf V$}).
@@
We can now see the primary motivation for the common literature assumptions (3) (assume dominant neutrals with ${\bf U}\approx{\bf u}_{n}$, and neglect all $\dot{\rho}_{ij}$ and non-neutral collision $\omega_{ij}$ terms) and (4) (assume vanishing drift/slip so $\omega_{ij}\rightarrow\omega_{ij}({\bf V}=\boldsymbol{0})$). With these simplifications, $\bf S$} has a trivial block form and the equation for each species is separable: ${\bf W}\equiv(c/B)\,({\bf E}+{\bf U}\times{\bf B}/c)\approx{\bf b}\times\delta{\bf u}_{jn}+\tilde{\beta}_{j}^{-1}\delta{\bf u}_{jn}$ (where $\delta{\bf u}_{jn}\equiv{\bf u}_{j}-{\bf u}_{n}$ are the slip velocities and $\tilde{\beta}_{j}\equiv q_{j}B/m_{j}c\omega_{jn}$). This gives the familiar result that the relative contributions to the current for each species depend only on $\tilde{\beta}_{j}$, and coefficients $\eta_{O}=(cB/4\pi)\tilde{\sigma}_{O}^{-1}$, $\eta_{H}=(cB/4\pi)\tilde{\sigma}_{H}/\tilde{\sigma}_{\bot}^{2}$, $\eta_{A}=(cB/4\pi)(\tilde{\sigma}_{P}/\tilde{\sigma}_{\bot}^{2}-1/\tilde{\sigma}_{O})$ where $\tilde{\sigma}_{O}\equiv\sum n_{j}q_{j}\tilde{\beta}_{j}$, $\tilde{\sigma}_{H}\equiv n_{j}q_{j}/(1+\tilde{\beta}_{j}^{2})$, $\tilde{\sigma}_{P}\equiv n_{j}q_{j}\tilde{\beta}_{j}/(1+\tilde{\beta}_{j}^{2})$, $\tilde{\sigma}_{\bot}^{2}\equiv\tilde{\sigma}_{H}^{2}+\tilde{\sigma}_{P}^{2}$. With the slip velocities [^7] The slip velocities are given in a frame where $^{\bf b}=\{0,0,1\}=\hat{z}$ by: $\delta{\bf u}_{jn,x}$ $=\frac{\tilde{\beta}_{j}(W_{x}+\tilde{\beta}_{j}W_{y})}{(1+\tilde{\beta}_{j}^{2})}=\frac{\tilde{\beta}_{j}[J_{x}(\tilde{\sigma}_{P}-\tilde{\beta}_{j}\tilde{\sigma}_{H})+J_{y}(\tilde{\sigma}_{H}+\tilde{\beta}_{j}\tilde{\sigma}_{P})]}{(1+\tilde{\beta}_{j}^{2})\sigma_{\bot}^{2}}\ ,$ $\delta{\bf u}_{jn,y}$ $=\frac{\tilde{\beta}_{j}(W_{y}-\tilde{\beta}_{j}W_{x})}{(1+\tilde{\beta}_{j}^{2})}=\frac{\tilde{\beta}_{j}[J_{y}(\tilde{\sigma}_{P}-\tilde{\beta}_{j}\tilde{\sigma}_{H})-J_{x}(\tilde{\sigma}_{H}+\tilde{\beta}_{j}\tilde{\sigma}_{P})]}{(1+\tilde{\beta}_{j}^{2})\sigma_{\bot}^{2}}\ ,$ $\delta{\bf u}_{jn,z}$ $=\tilde{\beta}_{j}W_{z}=\frac{\tilde{\beta}_{j}J_{z}}{\tilde{\sigma}_{O}}\ .$ (24) $\delta{\bf u}_{j}$, we can calculate each pairwise drift ($\delta{\bf u}_{j}-\delta{\bf u}_{i}$), and again in principle modify $\omega_{ij}$ to account for each drift+slip speed. But this again introduces the difficulties above: we then modify $\tilde{\beta}_{j}$ so must iteratively re-calculate the relative drift/slip speeds, are not ensured a unique solution, and (conceptually) must evaluate “which drifts matter.” [^8] For example, even if electrons carry most of the current, since we only retain their effective collision rate with neutrals here, do we include anomalous resistivity if their drift is subthermal relative to a dominant ion in the current but superthermal relative to a more trace ion, like charged grains?
@@
Figure 4: Ohmic, Hall, and ambipolar diffusivities $\eta_{O,H,A}$ as a function of density as in Fig. 2 (for the “circumstellar disk”-like case), but for the specific numerical examples in § A.3. Left: Comparison of a generally-subthermal case ($\ell_{B}=10\,\ell_{\rm crit}$; dotted) to a superthermal drift case ($\ell_{B}=0.1\,\ell_{\rm crit}^{\rm drift}$; solid). Right: Same but for the superthermal slip case $\ell_{B}=0.1\,\ell_{\rm crit}^{\rm slip}$. For each, we compare (1) the prediction from the simple parameterized rescaling we propose in Eq. 16 (for the multi-species scalings in § 2.4.2; “Approximate” in thin lines); to (2) the case where ignore charged dust grains as a current carrier (treat them as part of the neutral inertia taking $\tilde{\beta}_{\rm grain}\rightarrow 0$) and solve the full non-linear three-species problem with the exact coefficients (§ 2.1; “Exact” 3-species in intermediate lines); and (2) the results of exactly solving the non-linear multi-species equations as described in § A.1 with all possible pairs of species featuring simple Epstein-type scalings and anomalous terms (§ A.3; “Exact” 5-species in thick lines). The 3 and 5-species methods give excellent agreement, indicating that grain current and differences between exact velocity-dependent scalings in the superthermal regime make little difference. The closed-form analytic approximate scaling from Eq. 16 reproduces these reasonably well in all regimes of interest. Key qualitative results (e.g. vanishing of the Hall-dominated regime) are robust.
@@ A.2 Application to Scalings In the Text
A.3 Numerical Examples
To numerically validate the above and our approach, we revisit our simple example in the text of Fig. 2, where we calculate the $\eta_{O,H,A}$ coefficients according to some chemical model, but solve the generalized non-linear multidimensional problem exactly as described above (§ A.1; making assumptions (1), (2), (5) to close the system, but no others). This cannot be done in NICIL without extensive code modifications, so we instead adopt an extremely simple chemistry model for the quantities needed to compute the values shown. Specifically, we begin from (with all units here in CGS) similar parameters to those in the NICIL default model in-text (Wurster, 2021) with temperature $T=10\sqrt{1+(n/10^{n_{T,1}})^{0.8}}(1+n/10^{n_{T,2}})^{-0.3}(1+n/10^{n_{T,3}})^{0.57}$ for $(n_{T,1},n_{T,2},n_{T,3})=(11,16,21)$ from Machida et al. (2006); and $B=10^{-6}n^{1/2}$ for $n<n_{0,B}$, $B=10^{-6}(nn_{0,B})^{1/4}$ for $n\geq n_{0,B}$ with $n_{0,B}=10^{6}$ from Wardle & Ng (1999). We adopt a simple five-species model of: molecular neutrals with Solar abundances ($\xi_{n}\approx 1$; $m_{n}=2.3m_{p}$); electrons ($n_{e}\sim 0.01$ for $n<10^{n_{e,1}}$ and $n_{e}\sim 0.01(n/10^{n_{e,1}})^{2}$ for $n\geq 10^{n_{e,1}}$, with $n_{e,1}=14$); atomic+molecular ions ($n_{i}=n_{e}$, $q_{i}=e$, $\langle m_{i}\rangle=29m_{p}$); negative and positive small dust grains ($n_{g+}=n_{g-}\sim 10^{-13}n$, $q_{g+}=-q_{g-}=e$, $R_{\rm grain}=3\times 10^{-6}$). To be more appropriate for a circumstellar disk, we adjust the model parameters above to give similar results to the minimum mass Solar nebula (MMSN) model from Chiang & Youdin (2010), taking $(n_{T,1},n_{T,2},n_{T,3})\rightarrow(10.5,12.5,19.5)$, $n_{e,1}\rightarrow 20$, and $n_{g\pm}\rightarrow 10^{-15}n$ at $n<10^{n_{e,1}}$. We take the zero-drift cross sections to have values $\langle\sigma v\rangle_{in}=1.5\times 10^{-9}$, $\langle\sigma v\rangle_{en}=4.7\times 10^{-10}T^{1/2}$, $\langle\sigma v\rangle_{gn}=5.1\times 10^{-6}T^{1/2}$, $\langle\sigma v\rangle_{ei}=51T^{-3/2}$ (Spitzer & Härm, 1953) and neglect any other reactions.
@@ A.3 Numerical Examples
We now have interactions between electrons, heavy ions, neutrals, positively and negatively charged grains. Per § A.1, to even write down the equations for $\bf E$} in this limit we must define the velocity-dependent interaction terms between all pairs of particles, each of which can have different instability thresholds and qualitative behaviors of their anomalous terms (many of which remain poorly understood in detail). So we necessarily adopt a simple toy model, with the purpose not of providing any sort of “first principles” calculation, but simply to validate the simple approximate scalings discussed above. We neglect charge-exchange reactions, but consider all effective collisional cross terms. This includes both a “classical” (but correctly velocity-dependent) rate which we approximate with a simple hard-sphere-scattering type expression $\omega^{\rm coll}_{ji}=\omega^{\rm coll}_{ji}(\delta{\bf u}_{ji}\rightarrow 0)\times\left[1+|\delta{\bf u}_{ji}|^{2}/v_{T,\,ij}^{2}\right]^{1/2}$ (for which we use the values of $\langle\sigma v(\delta{\bf u}_{ji}\rightarrow 0)\rangle$ in text, supplemented with the zero-drift ion-grain and electron-grain cross sections from Draine & Sutin 1987), plus a pairwise anomalous term to capture different instabilities that can occur between each pair of interacting charged species, $\omega^{\rm an}_{ji}={\rm MAX}[\Omega_{j},\,\Omega_{i},\,\omega_{pj},\,\omega_{pi}]\,\Theta(|\delta{\bf u}_{ji}|/v_{T,ij})$ for $j,\,i\neq n$, with $v_{T,\,ij}={\rm MIN}[v_{T,\,i},\,v_{T,\,j}]$ (per Eqs. 20-22, with $\psi_{j}=0$ for $j\neq i,n$, so just the effective speeds of the individual carriers plus potentially-entrained neutrals for each). Like in Fig. 2 we consider two limits, which we define by the magnetic-field gradient $\ell_{B}=10\,\ell_{\rm crit}$ and $\ell_{B}=0.1\,\ell_{\rm crit}^{\rm drift}$ or $\ell_{B}=0.1\,\ell_{\rm crit}^{\rm slip}$ (with $v_{T}={\rm MIN}[v_{T,\,n},\,v_{A,\,{\rm ideal}}]$ in $\ell_{\rm crit}$). In the model in the text, these correspond one-to-one to the sub-thermal ($v_{\rm drift}\sim 0.1\,v_{T}$) and super-thermal ($v_{\rm drift}\sim 10\,v_{T}$) limits, but since there is no single drift velocity to speak of in the more general case, they must be defined in terms of $\ell_{B}$ (which is independent of the collision rates and species involved). Because of the non-linear dependencies involved, we must also specify the geometry, namely $\cos{\theta}\equiv\hat{\bf J}\cdot{\bf b}$, for which we take a representative $\theta=30^{\circ}$, but note the results are very weakly dependent on $\theta$, except if very close to parallel ($|\cos{\theta}|\approx 1$) where the ambipolar and Hall terms trivially vanish. Again, we emphasize the various coefficients and rates above are not meant to be exact, but simply to provide some calculable illustration.
@@
We further consider one more comparison. If we treat the dust grains as simply a “charge sink” and assume their mass to be anchored to the neutrals (adding to neutral inertia to leading order), then our system reduces to a three-component (neutrals, ions, electrons) system. We take this, plus the full velocity-dependent scattering rates from Pinto & Galli (2008) for each component, with a single ion-electron anomalous resistivity added to the ion-electron collision rate (where present), and otherwise adopt the scalings above. Then we can apply our solution from the main text (§ 2.1) and exactly solve the non-linear problem with all terms from § 2.1 included (still making assumptions (1), (2), (5) required for closure).